OpenStudy (anonymous):

A car accelerates at a steady rate from 0 to 108km per hour in 30 seconds, travels at a constant speed for one minute, and then brakes at a constant rate to stop in 10 seconds. i) Draw a velocity-time graph with velocity (vertical axis) in m/s and time (horizontal axis) in seconds. ii) Sketch a displacement-time graph for the motion.

OpenStudy (asnaseer):

 \(v_1\) on the graph represents 108km/h converted to m/s:\[v_1=\frac{108*1000}{60*60}=30\text{m/s}\] speed = distance/time, so: distance = speed * time which means the displacement-time graph is basically calculated by working out the area under the velocity-time graph.